10. Given that w denotes either one of the non-real roots of the equation z
3 = 1, show
that
(a) 1 + w + w
2 = 0,
1
Expert's answer
2020-05-20T19:17:34-0400

Suppose that "w" is one of the non-real roots of the equation "z^3=1" .

"w^3=1, \ \ w^3-1=0, \ \ (w-1)(1+w+w^2)=0 \ \"

"w \ \bcancel{=}\ 1," because 1 is real root of the equation "z^3=1" . So, "w-1 \ \bcancel{=}\ 0."

We have that "1+w+w^2=0."